OpenStudy (anonymous):

A student makes a $10.50 purchase at the bookstore with a $20 bill the store has no bills and gives the change in quarters and fifty-cent pieces there are 30coins in all how many of each kind are there?

OpenStudy (anonymous):

Harkirat I did not see the answer

OpenStudy (anonymous):

Pls clear for me 1$ = 100 cents quarter = 25 cents ?? Correct???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so i giv th sol now...

OpenStudy (anonymous):

the amount that the store has to return is $9.50 i.e 950 cents Let the number of quarters i.e. 25 cent coins be = x Then number of 50 cent coins = (30 -x) [since total no of coins is 30] Value of x coins of 25 cents = 25x cents value of (30-x) coins of 50 cents = 50 (30-x) = (1500 -50x) ATQ, 25x + (1500 -50x) = 950 -25x + 1500 = 950 -25x = 950 - 1500 x = -550/-25 x = 22 So number of quarter coins is 22 and number of 50 cent coins is (30-22) = 8 Clear???

OpenStudy (anonymous):

if clear, pls click GOOD ANSWER button